OpenStudy (anonymous):

I really thought someone could help me , man i was such a fool to think someone would help me. my friends told me about this site. but yea stupid me.

OpenStudy (anonymous):

NaCl

OpenStudy (anonymous):

post the question....

OpenStudy (anonymous):

you'll need to post ur question so everyone can see it.... here i copy/pasted for u: x+y=-2 x+y=3 use the method of substitution to solve the system of linear equation. these system may have, no solution, one solution, or infinitely many solution. honestly I do not know how to do this.

OpenStudy (anonymous):

in solving a system of linear equations, it is equivalent to finding where the two lines intersect. and just by the look of the two equations, notice the left side of the equations are exactly alike, but the right sides are different. this means the two lines are parallel and they don't intersect. therefore, there is no solution to this system....

OpenStudy (anonymous):

but since you asked about solving this using substitution, we will do that.... first, can you solve for a variable in the first equation.... solve for x.... ????

OpenStudy (anonymous):

do not do the problem by messaging.... you need to communicate with me (or any other helper) using the dialogue box like i'm doing here....

OpenStudy (anonymous):

okay, I was saying is it like y=-2 +y=3 =x-1

OpenStudy (anonymous):

solving that first equation for x, you get: x = -y-2 we can substitute this into the second equation: x + y = 3 (-y-2) + y = 3 -y - 2 + y = 3 -2 = 3 see the last line??? when is -2 = 3 ??? never... so there is no solution.

OpenStudy (anonymous):

be careful.. solving that first equation for x should give you -y - 2 NOT -y + 2....

OpenStudy (anonymous):

okay, I thought it would be x 1 . thank you so much. you help me a lot, I hope you didn't get upset with me and say oh my goodness this girl is messed up. thank you so much from the bottom of my heart for helping me. God Bless You

OpenStudy (anonymous):

yw... :) and i never did think that....